# A Dynamical Theory of Overfitting and Generalization in Large Two-Layer Networks

### v2

@@ Appendix I Lower bounding the overfitting timescale
where the ${\varepsilon}_{i}$ are i.i.d. Radamacher random variables and in the last two lines it is understood that the supremum is over $\|{\boldsymbol{w}}\|,\|\overline{\boldsymbol{w}}\|,\|{\boldsymbol{u}}\|\leq 1$. Consider the second term in the last expression. Defining $\eta(x)=x{\boldsymbol{1}}_{|x|\leq M}+M({\boldsymbol{1}}_{x>M}-{\boldsymbol{1}}_{x<-M})$, and $\overline{\eta}(x)=x-\eta(x)$, we have
$B_{2}$ $=2{\overline{a}}\,{\mathbb{E}}\sup_{{\boldsymbol{w}},\overline{\boldsymbol{w}},{\boldsymbol{u}}\in{\sf B}^{d}(1)}\frac{1}{n}\sum_{i=1}^{n}{\varepsilon}_{i}\sigma^{\prime}({\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i})\sigma(\overline{\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i}){\boldsymbol{u}}^{{\sf T}}{\boldsymbol{x}}_{i}$
$\leq 2{\overline{a}}\,{\mathbb{E}}\sup_{{\boldsymbol{w}},\overline{\boldsymbol{w}},{\boldsymbol{u}}\in{\sf B}^{d}(1)}\frac{1}{n}\sum_{i=1}^{n}{\varepsilon}_{i}\sigma^{\prime}({\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i})\sigma(\overline{\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i})\eta({\boldsymbol{u}}^{{\sf T}}{\boldsymbol{x}}_{i})$
$\phantom{AAAAA}+2{\overline{a}}\,{\mathbb{E}}\sup_{{\boldsymbol{w}},\overline{\boldsymbol{w}},{\boldsymbol{u}}\in{\sf B}^{d}(1)}\frac{1}{n}\sum_{i=1}^{n}{\varepsilon}_{i}\sigma^{\prime}({\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i})\sigma(\overline{\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i})\overline{\eta}({\boldsymbol{u}}^{{\sf T}}{\boldsymbol{x}}_{i})$
@@
$B_{2,1}={\overline{a}}\,{\mathbb{E}}\sup_{{\boldsymbol{w}},\overline{\boldsymbol{w}},{\boldsymbol{u}}\in{\sf B}^{d}(1)}\frac{1}{n}\sum_{i=1}^{n}{\varepsilon}_{i}\phi({\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i},\overline{\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i},{\boldsymbol{u}}^{{\sf T}}{\boldsymbol{x}}_{i})\,.$ (I.8)
@@
$B_{2,1}$ $\leq CL^{2}M{\overline{a}}\left\{{\mathbb{E}}\sup_{{\boldsymbol{w}}\in{\sf B}^{d}(1)}\frac{1}{n}\sum_{i=1}^{n}{\varepsilon}_{i}{\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i}+{\mathbb{E}}\sup_{\overline{\boldsymbol{w}}\in{\sf B}^{d}(1)}\frac{1}{n}\sum_{i=1}^{n}{\varepsilon}_{i}\overline{\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i}+{\mathbb{E}}\sup_{{\boldsymbol{u}}\in{\sf B}^{d}(1)}\frac{1}{n}\sum_{i=1}^{n}{\varepsilon}_{i}{\boldsymbol{u}}^{{\sf T}}{\boldsymbol{x}}_{i}\right\}$
@@
$\leq 2{\overline{a}}L^{2}\sup_{{\boldsymbol{u}}\in{\sf B}^{d}(1)}{\mathbb{E}}|\overline{\eta}({\boldsymbol{u}}^{{\sf T}}{\boldsymbol{x}}_{i})|+2{\overline{a}}L^{2}{\mathbb{E}}\sup_{{\boldsymbol{u}}\in{\sf B}^{d}(1)}\frac{1}{n}\sum_{i=1}^{n}{\varepsilon}_{i}|\overline{\eta}({\boldsymbol{u}}^{{\sf T}}{\boldsymbol{x}}_{i})|$
@@
The proof of Eq. (I.6) is completed by bounding $B_{1}$ along the same lines.
@@
$\leq 2\,{\mathbb{E}}\sup_{\boldsymbol{W},{\boldsymbol{u}}}\frac{1}{n}\sum_{i=1}^{n}{\varepsilon}_{i}y_{i}\sigma({\boldsymbol{w}}_{\ell}^{{\sf T}}{\boldsymbol{x}}_{i})+2\,{\mathbb{E}}\sup_{{\boldsymbol{a}},\boldsymbol{W},{\boldsymbol{u}}}\frac{1}{n}\sum_{i=1}^{n}{\varepsilon}_{i}\sum_{j=1}^{m}\frac{a_{j}}{m}\sigma({\boldsymbol{w}}_{j}^{{\sf T}}{\boldsymbol{x}}_{i})\sigma({\boldsymbol{w}}_{\ell}^{{\sf T}}{\boldsymbol{x}}_{i})$
@@
$D_{2}$ $\leq 2{\overline{a}}\,{\mathbb{E}}\sup_{\boldsymbol{W},{\boldsymbol{u}}}\max_{j\leq m}\frac{1}{n}\sum_{i=1}^{n}{\varepsilon}_{i}\sigma({\boldsymbol{w}}_{j}^{{\sf T}}{\boldsymbol{x}}_{i})\sigma({\boldsymbol{w}}_{\ell}^{{\sf T}}{\boldsymbol{x}}_{i})$
$\leq 2{\overline{a}}\,{\mathbb{E}}\sup_{{\boldsymbol{w}},\overline{\boldsymbol{w}}\in{\sf B}^{d}(1)}\frac{1}{n}\sum_{i=1}^{n}{\varepsilon}_{i}\psi({\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i},\overline{\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i})$
@@
Term $D_{1}$ is controlled analogously, yielding the proof of Eq. (I.7). ∎
@@
$\mathscrsfs{R}({\boldsymbol{a}},\boldsymbol{W})=\frac{1}{2}(\tau^{2}+\|\varphi\|^{2})-\frac{1}{m}\sum_{i=1}^{m}a_{i}\widehat{\varphi}({\boldsymbol{U}}^{{\sf T}}{\boldsymbol{w}}_{i})+\frac{1}{2m^{2}}\sum_{i,j=1}^{m}a_{i}a_{j}h({\boldsymbol{w}}_{i}^{{\sf T}}{\boldsymbol{w}}_{j})\,.$ (I.9)
Lemma I.2.
@@
$\big{\|}\nabla_{{\boldsymbol{w}}_{\ell}}\mathscrsfs{R}({\boldsymbol{a}},\tilde{\boldsymbol{W}})-\nabla_{{\boldsymbol{w}}_{\ell}}\mathscrsfs{R}({\boldsymbol{a}},\boldsymbol{W})\big{\|}$ $\leq CL^{2}\,\frac{|a_{\ell}|}{m}\,(1+{\overline{a}})\max_{j\leq m}\big{\|}\tilde{\boldsymbol{w}}_{j}-{\boldsymbol{w}}_{j}\big{\|}\,$, (I.10)
$\big{|}\partial_{a_{\ell}}\mathscrsfs{R}({\boldsymbol{a}},\tilde{\boldsymbol{W}})-\partial_{a_{\ell}}\mathscrsfs{R}({\boldsymbol{a}},\boldsymbol{W})\big{|}$ $\leq\frac{CL^{2}}{m}\,(1+{\overline{a}})\max_{j\leq m}\big{\|}\tilde{\boldsymbol{w}}_{j}-{\boldsymbol{w}}_{j}\big{\|}\,$, (I.11)
@@
$\big{\|}\nabla_{{\boldsymbol{w}}_{\ell}}\mathscrsfs{R}({\boldsymbol{\tilde{a}}},\boldsymbol{W})-\nabla_{{\boldsymbol{w}}_{\ell}}\mathscrsfs{R}({\boldsymbol{a}},\boldsymbol{W})\big{\|}$ $\leq\frac{CL^{2}}{m}(1+{\overline{a}})|\tilde{a}_{\ell}-a_{\ell}|+CL^{2}\frac{|a_{\ell}|}{m^{2}}\|{\boldsymbol{\tilde{a}}}-{\boldsymbol{a}}\|_{1}\,$, (I.12)
$\big{|}\partial_{a_{\ell}}\mathscrsfs{R}({\boldsymbol{\tilde{a}}},\boldsymbol{W})-\partial_{a_{\ell}}\mathscrsfs{R}({\boldsymbol{a}},\boldsymbol{W})\big{|}$ $\leq\frac{CL^{2}}{m^{2}}\,\big{\|}{\boldsymbol{\tilde{a}}}-{\boldsymbol{a}}\big{\|}_{1}\,.$ (I.13)
@@
Proof of Eq. (I.10). Differentiating Eq. (I.9)
$\frac{m}{a_{\ell}}\nabla_{{\boldsymbol{w}}_{\ell}}\mathscrsfs{R}({\boldsymbol{a}},\boldsymbol{W})=-{\boldsymbol{U}}\nabla\widehat{\varphi}({\boldsymbol{U}}^{{\sf T}}{\boldsymbol{w}}_{\ell})+\sum_{j=1}^{m}\frac{a_{j}}{m}h^{\prime}({\boldsymbol{w}}_{\ell}^{\sf T}{\boldsymbol{w}}_{j}){\boldsymbol{w}}_{j}\,.$ (I.14)
@@
Substituting above, this yields the claim (I.10).
Proof of Eq. (I.11). We proceed analogously to the previous point. Namely
$m\partial_{a_{\ell}}\mathscrsfs{R}({\boldsymbol{a}},\boldsymbol{W})=-\widehat{\varphi}({\boldsymbol{U}}^{{\sf T}}{\boldsymbol{w}}_{\ell})+\sum_{j=1}^{m}\frac{a_{j}}{m}h({\boldsymbol{w}}_{\ell}^{\sf T}{\boldsymbol{w}}_{j})\,$, (I.15)
@@
which implies immediately Eq. (I.11)
Proof of Eq. (I.12). Recalling Eq. (I.14), we have
@@
Proof of Eq. (I.12). Recalling Eq. (I.15), we have
@@
Corollary I.3.
Under the assumptions of Lemma I.2, there exists an absolute constant $C$ such that, for all $({\boldsymbol{a}},\boldsymbol{W}),({\boldsymbol{a}},\tilde{\boldsymbol{W}})\in{\mathcal{W}}_{m,d}^{\infty}({\overline{a}})$:
@@
$\begin{split}\dot{{\boldsymbol{a}}}_{0}(t)&=-m\nabla_{{\boldsymbol{a}}}\mathscrsfs{R}({\boldsymbol{a}}_{0}(t),\boldsymbol{W}_{0}(t))\,,\\ \dot{{\boldsymbol{w}}}_{0,i}(t)&=-m{\boldsymbol{P}}^{\perp}_{{\boldsymbol{w}}_{i}}\nabla_{{\boldsymbol{w}}_{i}}\mathscrsfs{R}({\boldsymbol{a}}_{0}(t),\boldsymbol{W}_{0}(t))\;\;\;\;\forall i=1,\ldots,m\,.\end{split}$ (I.17)
Lemma I.4.
@@
$T_{*}(m;c):=\inf\Big{\{}t:\,\big{(}\|{\boldsymbol{a}}(t)\|_{\infty}\vee\|{\boldsymbol{a}}_{0}(t)\|_{\infty}\big{)}\geq\Big{(}c_{*}L^{-2}\log\frac{ne}{d}\Big{)}^{1/3}\Big{\}}\wedge\Big{(}c_{*}L^{-2}\log\frac{ne}{d}\Big{)}^{1/3}\,.$ (I.18)
@@
$\sup_{t\leq T_{*}(m;c)}\Delta(t)\leq C(L^{2}+\tau^{2})\left(\frac{d}{n}\right)^{1/2-\delta}\,,\;\;\;\Delta(t):=\max_{\ell\leq m}\|\tilde{\boldsymbol{w}}_{\ell}(t)-{\boldsymbol{w}}_{\ell}(t)\|+\|{\boldsymbol{\tilde{a}}}(t)-{\boldsymbol{a}}(t)\|_{\infty}\,.$ (I.19)
@@
We will prove that the desired bound holds on the high-probability event of Lemma I.1, where by we set ${\overline{a}}=(c_{1}L^{-2}\log ne/d)^{1/3}$. Throughout the proof, we use $c_{0}$, $c_{1}$, $C$ to denote constants that might change from line to line, with dependence on the parameters of the problem as per the statement of the lemma.
@@
Since we are working on the event of Lemma I.1, and using Corollary I.3, we get, for $t\leq T_{*}(m;c)$.
@@
$\frac{{\rm d}\phantom{t}}{{\rm d}t}\|{\boldsymbol{w}}_{i}(t)-{\boldsymbol{w}}_{0,i}(t)\|\leq$ $C{\overline{a}}(L^{2}{\overline{a}}+\tau^{2})\sqrt{\frac{d}{n}\log(ne/d)}+CL^{2}(1+{\overline{a}}^{2})\Delta(t)\,.$ (I.20)
@@
Using the last bound together with Eq. (I.20), we get
@@
Lemma I.5 (Corollary 1 and Proposition 3 [BMZ24]).
Let ${\boldsymbol{a}}_{0}(t)$, $\boldsymbol{W}_{0}(t)$ be GF with respect to the population risk (I.17) with initialization $|a_{0,i}(0)|\leq a_{0}$ and $({\boldsymbol{w}}_{0,i}(0))_{i\leq m}\sim{\sf Unif}(S^{d-1})$. Recall that ${a}^{\mbox{\tiny\rm mf1}}_{i}(t)$, ${{\boldsymbol{v}}}^{\mbox{\tiny\rm mf1}}(t)$ is the solution of the ODEs (3.4) with initialization ${a}^{\mbox{\tiny\rm mf1}}_{i}(0)=a_{0,i}(0)$, ${{\boldsymbol{v}}}^{\mbox{\tiny\rm mf1}}_{i}(t)=0$. Under the assumptions of Theorem 3.2, for any ${\varepsilon}>0$ there exists constants $c_{0},c_{1}$ depending uniquely on $L$, and an absolute constant $C$ such that letting $T_{\mbox{\rm\tiny lb}}(m)=((c_{0}/{\varepsilon})\log m)^{1/3}$, the following happens with probability at least $1-2\exp(-c_{1}d)$,
$\sup_{t\leq T_{\mbox{\rm\tiny lb}}(m)}\frac{1}{m}\sum_{i=1}^{m}\Big{(}|a_{i}(t)-{a}^{\mbox{\tiny\rm mf1}}_{i}(t)|+\|{\boldsymbol{v}}_{i}(t)-{{\boldsymbol{v}}}^{\mbox{\tiny\rm mf1}}_{i}(t)\|\Big{)}$ $\leq C\,m^{{\varepsilon}}\Big{\{}\frac{1}{\sqrt{m}}+\frac{1}{\sqrt{d}}\Big{\}}\,$, (I.21)
$\sup_{t\leq T_{\mbox{\rm\tiny lb}}(m)}\Big{(}\mathscrsfs{R}({\boldsymbol{a}}(t),\boldsymbol{W}(t))-e_{\mbox{\rm\footnotesize ts}}(t)\Big{)}$ $\leq C\,m^{{\varepsilon}}\Big{\{}\frac{1}{\sqrt{m}}+\frac{1}{\sqrt{d}}\Big{\}}\,.$ (I.22)
@@
Throughout the proof $L,\tau,\alpha$ are assumed to be fixed, and constants $C,c_{0},\dots$ depend on them and can change from line to line. We will further work on the high probability events of Theorem 3.1, Lemma I.4, and Lemma I.5. By Theorem 3.1, for all $t\leq T_{\mbox{\rm\tiny lb}}(m)$ we have $\|{\boldsymbol{a}}(t)\|_{\infty}\leq c_{2}(\log 2m)^{1/3}$ (where the constant $c_{2}$ can be made sufficiently small, by eventually reducing $c_{1}$). An analogous of of Theorem 3.1 for the population risk implies $\|{\boldsymbol{a}}_{0}(t)\|_{\infty}\leq c_{2}(\log 2m)^{1/3}$ as well for all $t\leq T_{\mbox{\rm\tiny lb}}(m)$. Hence we can apply Lemma I.4 and Lemma I.5, which yields the claim. ∎
Appendix J Dynamical mean field theory for non-Gaussian model

### v3

@@ Appendix I Details about SGD simulations
where the $\xi_{i}$ are i.i.d. Radamacher random variables and in the last two lines it is understood that the supremum is over $\|{\boldsymbol{w}}\|,\|\overline{\boldsymbol{w}}\|,\|{\boldsymbol{u}}\|\leq 1$. Consider the second term in the last expression. Defining $\eta(x)=x{\boldsymbol{1}}_{|x|\leq M}+M({\boldsymbol{1}}_{x>M}-{\boldsymbol{1}}_{x<-M})$, and $\overline{\eta}(x)=x-\eta(x)$, we have
$B_{2}$ $=2{\overline{a}}\,{\mathbb{E}}\sup_{{\boldsymbol{w}},\overline{\boldsymbol{w}},{\boldsymbol{u}}\in{\sf B}^{d}(1)}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\sigma^{\prime}({\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i})\sigma(\overline{\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i}){\boldsymbol{u}}^{{\sf T}}{\boldsymbol{x}}_{i}$
$\leq 2{\overline{a}}\,{\mathbb{E}}\sup_{{\boldsymbol{w}},\overline{\boldsymbol{w}},{\boldsymbol{u}}\in{\sf B}^{d}(1)}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\sigma^{\prime}({\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i})\sigma(\overline{\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i})\eta({\boldsymbol{u}}^{{\sf T}}{\boldsymbol{x}}_{i})$
$\phantom{AAAAA}+2{\overline{a}}\,{\mathbb{E}}\sup_{{\boldsymbol{w}},\overline{\boldsymbol{w}},{\boldsymbol{u}}\in{\sf B}^{d}(1)}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\sigma^{\prime}({\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i})\sigma(\overline{\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i})\overline{\eta}({\boldsymbol{u}}^{{\sf T}}{\boldsymbol{x}}_{i})$
@@
$B_{2,1}={\overline{a}}\,{\mathbb{E}}\sup_{{\boldsymbol{w}},\overline{\boldsymbol{w}},{\boldsymbol{u}}\in{\sf B}^{d}(1)}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\phi({\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i},\overline{\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i},{\boldsymbol{u}}^{{\sf T}}{\boldsymbol{x}}_{i})\,.$ (J.9)
@@
$B_{2,1}$ $\leq CL^{2}M{\overline{a}}\left\{{\mathbb{E}}\sup_{{\boldsymbol{w}}\in{\sf B}^{d}(1)}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}{\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i}+{\mathbb{E}}\sup_{\overline{\boldsymbol{w}}\in{\sf B}^{d}(1)}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\overline{\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i}+{\mathbb{E}}\sup_{{\boldsymbol{u}}\in{\sf B}^{d}(1)}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}{\boldsymbol{u}}^{{\sf T}}{\boldsymbol{x}}_{i}\right\}$
@@
$\leq 2{\overline{a}}L^{2}\sup_{{\boldsymbol{u}}\in{\sf B}^{d}(1)}{\mathbb{E}}|\overline{\eta}({\boldsymbol{u}}^{{\sf T}}{\boldsymbol{x}}_{i})|+2{\overline{a}}L^{2}{\mathbb{E}}\sup_{{\boldsymbol{u}}\in{\sf B}^{d}(1)}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}|\overline{\eta}({\boldsymbol{u}}^{{\sf T}}{\boldsymbol{x}}_{i})|$
@@
The proof of Eq. (J.7) is completed by bounding $B_{1}$ along the same lines.
@@
$\leq 2\,{\mathbb{E}}\sup_{\boldsymbol{W},{\boldsymbol{u}}}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}y_{i}\sigma({\boldsymbol{w}}_{\ell}^{{\sf T}}{\boldsymbol{x}}_{i})+2\,{\mathbb{E}}\sup_{{\boldsymbol{a}},\boldsymbol{W},{\boldsymbol{u}}}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\sum_{j=1}^{m}\frac{a_{j}}{m}\sigma({\boldsymbol{w}}_{j}^{{\sf T}}{\boldsymbol{x}}_{i})\sigma({\boldsymbol{w}}_{\ell}^{{\sf T}}{\boldsymbol{x}}_{i})$
@@
$D_{2}$ $\leq 2{\overline{a}}\,{\mathbb{E}}\sup_{\boldsymbol{W},{\boldsymbol{u}}}\max_{j\leq m}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\sigma({\boldsymbol{w}}_{j}^{{\sf T}}{\boldsymbol{x}}_{i})\sigma({\boldsymbol{w}}_{\ell}^{{\sf T}}{\boldsymbol{x}}_{i})$
$\leq 2{\overline{a}}\,{\mathbb{E}}\sup_{{\boldsymbol{w}},\overline{\boldsymbol{w}}\in{\sf B}^{d}(1)}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\psi({\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i},\overline{\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i})$
@@
Term $D_{1}$ is controlled analogously, yielding the proof of Eq. (J.8). ∎
@@
$\mathscrsfs{R}({\boldsymbol{a}},\boldsymbol{W})=\frac{1}{2}(\tau^{2}+\|\varphi\|^{2})-\frac{1}{m}\sum_{i=1}^{m}a_{i}\widehat{\varphi}({\boldsymbol{U}}^{{\sf T}}{\boldsymbol{w}}_{i})+\frac{1}{2m^{2}}\sum_{i,j=1}^{m}a_{i}a_{j}h({\boldsymbol{w}}_{i}^{{\sf T}}{\boldsymbol{w}}_{j})\,.$ (J.10)
Lemma J.2.
@@
$\big\|\nabla_{{\boldsymbol{w}}_{\ell}}\mathscrsfs{R}({\boldsymbol{a}},\tilde{\boldsymbol{W}})-\nabla_{{\boldsymbol{w}}_{\ell}}\mathscrsfs{R}({\boldsymbol{a}},\boldsymbol{W})\big\|$ $\leq CL^{2}\,\frac{|a_{\ell}|}{m}\,(1+{\overline{a}})\max_{j\leq m}\big\|\tilde{\boldsymbol{w}}_{j}-{\boldsymbol{w}}_{j}\big\|\,$, (J.11)
$\big|\partial_{a_{\ell}}\mathscrsfs{R}({\boldsymbol{a}},\tilde{\boldsymbol{W}})-\partial_{a_{\ell}}\mathscrsfs{R}({\boldsymbol{a}},\boldsymbol{W})\big|$ $\leq\frac{CL^{2}}{m}\,(1+{\overline{a}})\max_{j\leq m}\big\|\tilde{\boldsymbol{w}}_{j}-{\boldsymbol{w}}_{j}\big\|\,$, (J.12)
@@
$\big\|\nabla_{{\boldsymbol{w}}_{\ell}}\mathscrsfs{R}({\boldsymbol{\tilde{a}}},\boldsymbol{W})-\nabla_{{\boldsymbol{w}}_{\ell}}\mathscrsfs{R}({\boldsymbol{a}},\boldsymbol{W})\big\|$ $\leq\frac{CL^{2}}{m}(1+{\overline{a}})|\tilde{a}_{\ell}-a_{\ell}|+CL^{2}\frac{|a_{\ell}|}{m^{2}}\|{\boldsymbol{\tilde{a}}}-{\boldsymbol{a}}\|_{1}\,$, (J.13)
$\big|\partial_{a_{\ell}}\mathscrsfs{R}({\boldsymbol{\tilde{a}}},\boldsymbol{W})-\partial_{a_{\ell}}\mathscrsfs{R}({\boldsymbol{a}},\boldsymbol{W})\big|$ $\leq\frac{CL^{2}}{m^{2}}\,\big\|{\boldsymbol{\tilde{a}}}-{\boldsymbol{a}}\big\|_{1}\,.$ (J.14)
@@
Proof of Eq. (J.11). Differentiating Eq. (J.10)
$\frac{m}{a_{\ell}}\nabla_{{\boldsymbol{w}}_{\ell}}\mathscrsfs{R}({\boldsymbol{a}},\boldsymbol{W})=-{\boldsymbol{U}}\nabla\widehat{\varphi}({\boldsymbol{U}}^{{\sf T}}{\boldsymbol{w}}_{\ell})+\sum_{j=1}^{m}\frac{a_{j}}{m}h^{\prime}({\boldsymbol{w}}_{\ell}^{\sf T}{\boldsymbol{w}}_{j}){\boldsymbol{w}}_{j}\,.$ (J.15)
@@
Substituting above, this yields the claim (J.11).
Proof of Eq. (J.12). We proceed analogously to the previous point. Namely
$m\partial_{a_{\ell}}\mathscrsfs{R}({\boldsymbol{a}},\boldsymbol{W})=-\widehat{\varphi}({\boldsymbol{U}}^{{\sf T}}{\boldsymbol{w}}_{\ell})+\sum_{j=1}^{m}\frac{a_{j}}{m}h({\boldsymbol{w}}_{\ell}^{\sf T}{\boldsymbol{w}}_{j})\,$, (J.16)
@@
which implies immediately Eq. (J.12)
Proof of Eq. (J.13). Recalling Eq. (J.15), we have
@@
Proof of Eq. (J.13). Recalling Eq. (J.16), we have
@@
Corollary J.3.
Under the assumptions of Lemma J.2, there exists an absolute constant $C$ such that, for all $({\boldsymbol{a}},\boldsymbol{W}),({\boldsymbol{a}},\tilde{\boldsymbol{W}})\in{\mathcal{W}}_{m,d}^{\infty}({\overline{a}})$:
@@
$\begin{split}\dot{{\boldsymbol{a}}}_{0}(t)&=-m\nabla_{{\boldsymbol{a}}}\mathscrsfs{R}({\boldsymbol{a}}_{0}(t),\boldsymbol{W}_{0}(t))\,,\\ \dot{{\boldsymbol{w}}}_{0,i}(t)&=-m{\boldsymbol{P}}^{\perp}_{{\boldsymbol{w}}_{i}}\nabla_{{\boldsymbol{w}}_{i}}\mathscrsfs{R}({\boldsymbol{a}}_{0}(t),\boldsymbol{W}_{0}(t))\;\;\;\;\forall i=1,\ldots,m\,.\end{split}$ (J.18)
Lemma J.4.
@@
$T_{*}(m;c):=\inf\Big\{t:\,\big(\|{\boldsymbol{a}}(t)\|_{\infty}\vee\|{\boldsymbol{a}}_{0}(t)\|_{\infty}\big)\geq\Big(c_{*}L^{-2}\log\frac{ne}{d}\Big)^{1/3}\Big\}\wedge\Big(c_{*}L^{-2}\log\frac{ne}{d}\Big)^{1/3}\,.$ (J.19)
@@
$\sup_{t\leq T_{*}(m;c)}\Delta(t)\leq C(L^{2}+\tau^{2})\left(\frac{d}{n}\right)^{1/2-\delta}\,,\;\;\;\Delta(t):=\max_{\ell\leq m}\|\tilde{\boldsymbol{w}}_{\ell}(t)-{\boldsymbol{w}}_{\ell}(t)\|+\|{\boldsymbol{\tilde{a}}}(t)-{\boldsymbol{a}}(t)\|_{\infty}\,.$ (J.20)
@@
We will prove that the desired bound holds on the high-probability event of Lemma J.1, where by we set ${\overline{a}}=(c_{1}L^{-2}\log ne/d)^{1/3}$. Throughout the proof, we use $c_{0}$, $c_{1}$, $C$ to denote constants that might change from line to line, with dependence on the parameters of the problem as per the statement of the lemma.
@@
Since we are working on the event of Lemma J.1, and using Corollary J.3, we get, for $t\leq T_{*}(m;c)$.
@@
$\frac{{\rm d}\phantom{t}}{{\rm d}t}\|{\boldsymbol{w}}_{i}(t)-{\boldsymbol{w}}_{0,i}(t)\|\leq$ $C{\overline{a}}(L^{2}{\overline{a}}+\tau^{2})\sqrt{\frac{d}{n}\log(ne/d)}+CL^{2}(1+{\overline{a}}^{2})\Delta(t)\,.$ (J.21)
@@
Using the last bound together with Eq. (J.21), we get
@@
Lemma J.5 (Corollary 1 and Proposition 3 [BMZ24]).
Let ${\boldsymbol{a}}_{0}(t)$, $\boldsymbol{W}_{0}(t)$ be GF with respect to the population risk (J.18) with initialization $|a_{0,i}(0)|\leq a_{0}$ and $({\boldsymbol{w}}_{0,i}(0))_{i\leq m}\sim{\sf Unif}(S^{d-1})$. Recall that ${a}^{\mbox{\tiny\rm mf1}}_{i}(t)$, ${{\boldsymbol{v}}}^{\mbox{\tiny\rm mf1}}(t)$ is the solution of the ODEs (3.4) with initialization ${a}^{\mbox{\tiny\rm mf1}}_{i}(0)=a_{0,i}(0)$, ${{\boldsymbol{v}}}^{\mbox{\tiny\rm mf1}}_{i}(t)=0$. Under the assumptions of Theorem 3.2, for any ${\varepsilon}>0$ there exists constants $c_{0},c_{1}$ depending uniquely on $L$, and an absolute constant $C$ such that letting $T_{\mbox{\rm\tiny lb}}(m)=((c_{0}/{\varepsilon})\log m)^{1/3}$, the following happens with probability at least $1-2\exp(-c_{1}d)$,
$\sup_{t\leq T_{\mbox{\rm\tiny lb}}(m)}\frac{1}{m}\sum_{i=1}^{m}\Big(|a_{i}(t)-{a}^{\mbox{\tiny\rm mf1}}_{i}(t)|+\|{\boldsymbol{v}}_{i}(t)-{{\boldsymbol{v}}}^{\mbox{\tiny\rm mf1}}_{i}(t)\|\Big)$ $\leq C\,m^{{\varepsilon}}\Big\{\frac{1}{\sqrt{m}}+\frac{1}{\sqrt{d}}\Big\}\,$, (J.22)
$\sup_{t\leq T_{\mbox{\rm\tiny lb}}(m)}\Big(\mathscrsfs{R}({\boldsymbol{a}}(t),\boldsymbol{W}(t))-e_{\mbox{\rm\footnotesize ts}}(t)\Big)$ $\leq C\,m^{{\varepsilon}}\Big\{\frac{1}{\sqrt{m}}+\frac{1}{\sqrt{d}}\Big\}\,.$ (J.23)
@@
Throughout the proof $L,\tau,\alpha$ are assumed to be fixed, and constants $C,c_{0},\dots$ depend on them and can change from line to line. We will further work on the high probability events of Theorem 3.1, Lemma J.4, and Lemma J.5. By Theorem 3.1, for all $t\leq T_{\mbox{\rm\tiny lb}}(m)$ we have $\|{\boldsymbol{a}}(t)\|_{\infty}\leq c_{2}(\log 2m)^{1/3}$ (where the constant $c_{2}$ can be made sufficiently small, by eventually reducing $c_{1}$). An analogous of of Theorem 3.1 for the population risk implies $\|{\boldsymbol{a}}_{0}(t)\|_{\infty}\leq c_{2}(\log 2m)^{1/3}$ as well for all $t\leq T_{\mbox{\rm\tiny lb}}(m)$. Hence we can apply Lemma J.4 and Lemma J.5, which yields the claim. ∎
Appendix K Dynamical mean field theory for non-Gaussian model